# Robust Signal Detection with Quadratically Convex Orthosymmetric Constraints

### v2

@@ Appendix B Deferred Proofs of the Lemmas and Main Theorems
$\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}\widetilde{X}_{i}\right\rVert_{2}^{2}=\underbrace{\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}(\widetilde{X}_{i}-\mu)\right\rVert_{2}^{2}}_{:=\text{I}}+\underbrace{\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}}_{:=\text{II}}+2\underbrace{\left|S^{\prime}\right|\sum\limits_{i\in S^{\prime}}\left[P^{\star}(\widetilde{X}_{i}-\mu)\right]^{\top}\mu}_{:=\text{III}}.$ (B.10)
@@
For $\text{I}_{2,S^{\prime}}$, since we allow $S^{\prime}$ to be any subset of $S$, refined analysis is needed. Using the Hanson-Wright inequality A.10 and observing that $\left\lVert P^{\star}\right\rVert_{F}=\sqrt{k},\left\lVert P^{\star}\right\rVert_{2}\leq 1$, for each fixed $S^{\prime}\subset S$, we have
@@
Table 1: terms from (B.10) and corresponding means and variations.
@@
To clarify the structure of this section and highlight how the definitions, lemmas, and algorithms fit together, we provide the following roadmap. We depart from the definition of regularity, and arrive at the final complete testing procedure, Algorithm 8. The pseudo code is provided in Appendix C.
@@
Definition B.7 ($\omega$-regularity).
@@
B.7 immediately recovers B.6 if $\omega=\mathbf{1}_{S}$.
@@
$\beta_{1}=\epsilon N\left(\sqrt{k^{\star}}+\sqrt{N}\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\sqrt{N}\left\lVert\nu\right\rVert_{2}^{2},$ (B.11)
@@
$\left|\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\left[\epsilon N\sqrt{(k^{\star}+\left\lVert\nu\right\rVert_{2}^{2})\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\left\lVert\nu\right\rVert_{2}^{2}\right]$ (B.12)
@@
By Lemma A.18, we have at most $\left(\frac{e}{\epsilon}\right)^{\epsilon N}$ different possible $S$ in total. Setting the probability to be $\frac{\alpha}{6}\left(\frac{\epsilon}{e}\right)^{\epsilon N}$ and solving for $t$, we obtain
@@
Combining the results together, and using the union bound over all possible $T$, we have
@@
with probability greater than $1-\frac{\alpha}{3}$. The RHS is exactly $\beta_{2}$ in (B.11).
Check of (iii). Finally, we are going to evaluate the term $\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in S}Y_{j}\right\rangle$. Since this is equal to $\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}+\sum\limits_{i\in S}\left\langle Y_{i},\sum\limits_{j\in S\backslash i}Y_{j}\right\rangle$. We have obtained the bound for $\left\lVert Y_{i}\right\rVert_{2}^{2}$ in (i). For the term of inner product, we have
$\left\langle Y_{i},\sum\limits_{j\in S\backslash i}Y_{j}\right\rangle=\left\langle Y_{i}-\nu,\sum\limits_{j\in S\backslash i}(Y_{j}-\nu)\right\rangle+\left\langle Y_{i}-\nu,(N-1)\nu\right\rangle+\left\langle\nu,\sum\limits_{j\in S\backslash i}(Y_{j}-\nu)\right\rangle+(N-1)\left\lVert\nu\right\rVert_{2}^{2}.$
@@
Comparing with (B.12) and Definition B.6, we find that the required $\beta_{1}$ is
@@
which is exactly $\beta_{1}$ in (B.11). The proof is completed. ∎
By Lemma B.8, we should be able to control the inner product between the sum of samples on two sets $S$ and $S^{\prime}$ by the polarization identity. The next corollary exactly states this benefit.
Corollary B.9.
@@
Since each of $\S,S\SS\cup S^{\prime},S\cap S^{\prime},S\backslash$ $^{\prime},S^{\prime}\backslash$ S has cardinality no more than $2\epsilon N$ by assumption, by Definition B.6 and Lemma B.8, we know
@@
With Lemma A.19, it is not hard to extend the result in Corollary B.9 to the case dealing with the convex combination of the indicator vectors.
Corollary B.10.
@@
In Corollary B.9, there is no corresponding (ii) since that is exactly the third definition of regularity.
@@
We proof by direct calculation. By Lemma A.19, the vectors $a$ and $b$ can be represent as the convex combination of some indicator vectors, respectively. Let $a=\sum\limits_{k=1}^{M_{a}}\xi_{k}\mathbf{1}_{S_{k}},b=\sum\limits_{l=1}^{M_{b}}\eta_{l}\mathbf{1}_{S_{l}}$, where $S_{k},S_{l}\subset[N],1\leq k\leq N,1\leq l\leq N$. We have:
@@
Here in the equation, (1) is from Corollary B.9, and $R_{k,l}\leq C\beta_{2}$ for some constant $C$. Now that by Lemma A.19, $0\leq\sum\limits_{k=1}^{M_{a}}\xi_{k}\leq 1,0\leq\sum\limits_{l=1}^{M_{b}}\eta_{l}\leq 1$, we know $\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}R_{k,l}\leq C\beta_{2}$. Therefore, (i) is checked. With very similar argument, we can also check (ii). The proof is completed. ∎
At this stage, we have all tools needed for the proof. The goal of the following Algorithms 3, 4, 5 is to filter the weight and samples such that the resulting weight and samples meet the requirements of $\omega$-regularity. We start with the first requirement, which is the simplest one.
@@ 9 $\texttt{i}=\texttt{i}+1$
Lemma B.11.
Upon finishing the algorithm 3, all of samples satisfy
@@
Moreover, under the null hypothesis, the algorithm will terminate without rejecting with probability at least $1-\alpha$.
@@
The satisfaction of the inequality comes directly from the definition of Algorithm 3. According to Lemma B.8 where setting $\left\lVert\nu\right\rVert_{2}=0$, all uncontaminated samples will satisfy the inequality with probability at least $1-\alpha$ and therefore will not be deleted. The remaining follows. ∎
@@
$\gamma_{2}:=c\left[\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right].$ (B.13)
@@
1Set $\gamma_{2}$ as (B.13), and $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
@@ 11  Set $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$
Proof of (i) is relatively easy. By the update (B.14), the $\omega_{i}$ corresponding to the maximal $\tau_{i}$ will be set to zero in each iteration. Consequently, $\left\lVert\omega\right\rVert_{1}$ will be less than $N(1-2\epsilon)$ in at most $2\epsilon N$ rounds, followed by the rejection.
@@
Now it remains to prove the statement. Suppose $\omega^{(t)}\in\Lambda_{1}$ and we are still in the iteration. By the definition of $\Lambda_{1}$ and the update (B.14), it suffices to prove $\sum\limits_{i\in[I]\cap C}\tau_{i}^{(t)}\omega_{i}^{(t)}\geq\sum\limits_{i\in[I]\cap([N]\backslash C)}\tau_{i}^{(t)}\omega^{(t)}$, i.e.,
@@
By Corollary A.14 (set $\left\lVert\nu\right\rVert_{2}=0$) and the definition of $\gamma_{2}$ (B.13), with high probability, we have
@@
By the definition of $I$ we know $2\epsilon N\leq\sum\limits_{i\in[I]}\omega_{i}<2\epsilon N+1$. Because $\omega^{(t)}\in\Lambda_{1}$, we know $\left\lVert\mathbf{1}-\omega^{(t)}\right\rVert_{1}\leq 2\left\lVert\mathbf{1}_{C}-\omega_{C}^{(t)}\right\rVert_{1}\leq 2\epsilon N$. Hence $I-(2\epsilon N+1)\leq\sum\limits_{i\in[I]}(1-\omega_{i}^{(t)})\leq 2\epsilon N$, which leads to $I\leq 4\epsilon N+1$. Therefore, according to Lemma A.15, with high probability, we have
$v^{\top}\left(\sum\limits_{i\in[I]\cap([N]\backslash C)}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}\right)v\leq c\gamma_{2}.$ (B.15)
@@
Therefore $\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{2}{3}\sum\limits_{i\in C}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{2(1-c)}{3}\gamma_{2}$. Note that by selecting the constant in (B.13) sufficiently large, we can let $c$ sufficiently small. Hence, we can assume $c\leq\frac{1}{4}$, by which we have $\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{1}{2}\gamma_{2}$. Combine with (B.15), we find
@@
where the last step is from the definition of $\gamma_{2}$ (B.13). If we expand $v^{\top}Y_{i}$, this merely says
$\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]^{2}+\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}+2\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]\left\lVert\nu\right\rVert_{2}\leq cN.$ (B.16)
@@
Adding both sides with $\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}$ and plugging (B.16) in, we have
@@
Now since $\left\lVert v\right\rVert_{2}=1$, we know that $\left[v^{\top}(Y_{i}-\nu)\right]^{2}$ is subject to a chi-square distribution with degree of freedom 1. By the standard argument for the sub-exponential random variables (see A.7), we know that $\sum\limits_{i\in[N]\backslash C}\left[v^{\top}(Y_{i}-\nu)\right]^{2}\leq N+c\sqrt{N}$ with (high) constant probability. On the other hand, by the definition of the algorithm we have $\left\lVert\omega_{[N]\backslash C}\right\rVert_{1}\geq\left\lVert\omega\right\rVert_{1}-\epsilon N\geq(1-3\epsilon)N$. Therefore, we have with (high) constant probability,
@@
Now deviding both sides by $\frac{1}{2}(1-3\epsilon)N$, we see that Lemma B.13 holds if we take $c=\frac{2(c_{1}+c_{2})}{1-3\epsilon}$
We are going to prove (iii) by contradiction. Suppose that at the end of the algorithm without rejection, we have $\left\lVert\sum\limits_{i\in S}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}>2\gamma_{2}\epsilon N$ for some $S\subset[N]$ with $\left|S\right|\leq\epsilon N$. This indicates that there exists a vector $v\in\mathbb{R}^{k^{\star}}$ with $\left\lVert v\right\rVert_{2}=1$ such that $\sum\limits_{i\in S}\sqrt{\omega_{i}}v^{\top}Y_{i}>\sqrt{2\gamma_{2}\epsilon N}$. Therefore, by the Cauchy-Schwarz inequality, we have $\sum\limits_{i\in S\cup C}\omega_{i}(v^{\top}Y_{i})^{2}\geq 2\gamma_{2}$. This implies that
$\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}+\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}\right\rVert_{2}\geq\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}\right\rVert_{2}\geq 2\gamma_{2}.$ (B.17)
@@
$\leq\underbrace{\left\lVert\sum\limits_{i=1}^{N}\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}-N\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}}_{\text{I}}+\underbrace{\left\lVert\sum\limits_{i=1}^{N}(1-\omega_{i})\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}\right\rVert_{2}}_{\text{II}}$
$+\underbrace{\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}\widetilde{Y_{i}}\widetilde{Y_{i}}^{\top}\right\rVert}_{\text{III}}+\left|S\cup C\right|\left\lVert\nu\right\rVert_{2}^{2}.$
By Lemma B.13, without loss of generality we can assume $\left\lVert\nu\right\rVert_{2}\leq c_{0}$ for some constant $c_{0}$. For the last term, since $\left|S\cup C\right|\leq 2\epsilon N$, obviously it can be bounded by $c\gamma_{2}$. We are going to deal with the other three parts respectively.

### v3

@@ Appendix C Deferred Proofs of the Lemmas and Main Theorems
$\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}\widetilde{X}_{i}\right\rVert_{2}^{2}=\underbrace{\left\lVert P^{\star}\sum\limits_{i\in S^{\prime}}(\widetilde{X}_{i}-\mu)\right\rVert_{2}^{2}}_{:=\text{I}}+\underbrace{\left|S^{\prime}\right|^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}}_{:=\text{II}}+2\underbrace{\left|S^{\prime}\right|\sum\limits_{i\in S^{\prime}}\left[P^{\star}(\widetilde{X}_{i}-\mu)\right]^{\top}\mu}_{:=\text{III}}.$ (C.10)
@@
For $\text{I}_{2,S^{\prime}}$, since we allow $S^{\prime}$ to be any subset of $[N]$, refined analysis is needed. Using the Hanson–Wright inequality B.10 and observing that $\left\lVert P^{\star}\right\rVert_{F}=\sqrt{k^{\star}},\left\lVert P^{\star}\right\rVert_{2}\leq 1$, for each fixed $S^{\prime}\subset[N]$, we have
@@
Table 1: corresponding means and variations of the terms from (C.10).
@@
To clarify the structure of this section and highlight how the definitions, lemmas, and algorithms fit together, we provide the following roadmap. We depart from the definition of regularity, and arrive at the final complete testing procedure, Algorithm 8. The pseudo code is provided in Appendix A.
@@
Definition C.7 ($\omega$-regularity).
@@
C.7 immediately recovers C.6 if $\omega=\mathbf{1}_{S}$.
@@
$\beta_{1}=\epsilon N\left(\sqrt{k^{\star}}+\sqrt{N}\left\lVert\nu\right\rVert_{2}\right)\sqrt{\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\sqrt{N}\left\lVert\nu\right\rVert_{2}^{2},$ (C.11)
@@
$\left|\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}-\left|S\right|k^{\star}\right|\leq c\left[\epsilon N\sqrt{(k^{\star}+\left\lVert\nu\right\rVert_{2}^{2})\ln\left(\frac{N}{\alpha}\right)}+\epsilon N\ln\left(\frac{N}{\alpha}\right)+\epsilon N\left\lVert\nu\right\rVert_{2}^{2}\right]$ (C.12)
@@
By Lemma B.18, we have at most $\left(\frac{e}{\epsilon}\right)^{\epsilon N}$ different possible $S$ in total. Setting the probability to be $\frac{\alpha}{6}\left(\frac{\epsilon}{e}\right)^{\epsilon N}$ and solving for $t$, we obtain
@@
Combining the results together, and using the union bound over all possible $S$, we have
@@
with probability greater than $1-\frac{\alpha}{3}$. The RHS is exactly $\beta_{2}$ in (C.11).
Check of (iii). Finally, we are going to evaluate the term $\left\langle\sum\limits_{i\in S}Y_{i},\sum\limits_{j\in[N]}Y_{j}\right\rangle$. Since this is equal to $\sum\limits_{i\in S}\left\lVert Y_{i}\right\rVert_{2}^{2}+\sum\limits_{i\in S}\left\langle Y_{i},\sum\limits_{j\in[N]\backslash i}Y_{j}\right\rangle$. We have obtained the bound for $\left\lVert Y_{i}\right\rVert_{2}^{2}$ in (i). For the term of inner product, we have
$\left\langle Y_{i},\sum\limits_{j\in[N]\backslash i}Y_{j}\right\rangle=\left\langle Y_{i}-\nu,\sum\limits_{j\in[N]\backslash i}(Y_{j}-\nu)\right\rangle+\left\langle Y_{i}-\nu,(N-1)\nu\right\rangle+\left\langle\nu,\sum\limits_{j\in[N]\backslash i}(Y_{j}-\nu)\right\rangle+(N-1)\left\lVert\nu\right\rVert_{2}^{2}.$
@@
Comparing with (C.12) and Definition C.6, we find that the required $\beta_{1}$ is
@@
which is exactly $\beta_{1}$ in (C.11). The proof is completed. ∎
By Lemma C.8, we should be able to control the inner product between the sum of samples on two sets $S$ and $S^{\prime}$ by the polarization identity. The next corollary exactly states this benefit.
Corollary C.9.
@@
Since each of $\S,S\SS\cup S^{\prime},S\cap S^{\prime},S\backslash$ $^{\prime},S^{\prime}\backslash$ S has cardinality no more than $2\epsilon N$ by assumption, by Definition C.6 and Lemma C.8, we know
@@
With Lemma B.19, it is not hard to extend the result in Corollary C.9 to the case dealing with the convex combination of the indicator vectors.
Corollary C.10.
@@
In Corollary C.9, there is no corresponding (ii) since that is exactly the third definition of regularity.
@@
We proof by direct calculation. By Lemma B.19, the vectors $a$ and $b$ can be represent as the convex combination of some indicator vectors, respectively. Let $a=\sum\limits_{k=1}^{M_{a}}\xi_{k}\mathbf{1}_{S_{k}},b=\sum\limits_{l=1}^{M_{b}}\eta_{l}\mathbf{1}_{S_{l}}$, where $S_{k},S_{l}\subset[N],1\leq k\leq N,1\leq l\leq N$. We have:
@@
Here in the equation, (1) is from Corollary C.9, and $R_{k,l}\leq c\beta_{2}$ for some constant $c$. Now that by Lemma B.19, $0\leq\sum\limits_{k=1}^{M_{a}}\xi_{k}\leq 1,0\leq\sum\limits_{l=1}^{M_{b}}\eta_{l}\leq 1$, we know $\sum\limits_{k=1}^{M_{a}}\sum\limits_{l=1}^{M_{b}}\xi_{k}\eta_{l}R_{k,l}\leq c\beta_{2}$. Therefore, (i) is checked. With very similar argument, we can also check (ii). The proof is completed. ∎
At this stage, we have all tools needed for the proof. The goal of the following Algorithms 9, 10, 11 is to filter the weight $\omega$ and the samples $\mathbf{Y}$ such that the resulting $\omega$ and $\mathbf{Y}$ meet the requirements of $\omega$-regularity. We start with the first requirement.
@@ 9 $\texttt{i}=\texttt{i}+1$
Lemma C.11.
Upon finishing Algorithm 9, all of samples satisfy
@@
Moreover, under the null hypothesis $H_{0}$, the algorithm will terminate without rejecting with probability at least $1-\alpha$.
@@
The satisfaction of the inequality comes directly from the definition of Algorithm 9. According to Lemma C.8 where setting $\left\lVert\nu\right\rVert_{2}=0$, all uncontaminated samples will satisfy the inequality with probability at least $1-\alpha$ and therefore will not be deleted. The remaining follows. ∎
@@
$\gamma_{2}:=c\left[\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right].$ (C.13)
@@
1Set $\gamma_{2}$ as (C.13), and $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
@@ 11  Set $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$
Proof of (i) is relatively easy. By the update (C.14), the $\omega_{i}$ corresponding to the maximal $\tau_{i}$ will be set to zero in each iteration. Consequently, $\left\lVert\omega\right\rVert_{1}$ will be less than $N(1-2\epsilon)$ in at most $2\epsilon N$ rounds, followed by the rejection.
@@
Now it remains to prove the statement. Suppose $\omega^{(t)}\in\Lambda_{1}$ and we are still in the iteration. By the definition of $\Lambda_{1}$ and the update (C.14), it suffices to prove $\sum\limits_{i\in[I]\cap C}\tau_{i}^{(t)}\omega_{i}^{(t)}\geq\sum\limits_{i\in[I]\cap([N]\backslash C)}\tau_{i}^{(t)}\omega^{(t)}$, i.e.,
@@
By Corollary B.14 (set $\left\lVert\nu\right\rVert_{2}=0$) and the definition of $\gamma_{2}$ (C.13), with high probability, we have
@@
By the definition of $I$ we know $2\epsilon N\leq\sum\limits_{i\in[I]}\omega_{i}<2\epsilon N+1$. Because $\omega^{(t)}\in\Lambda_{1}$, we know $\left\lVert\mathbf{1}-\omega^{(t)}\right\rVert_{1}\leq 2\left\lVert\mathbf{1}_{C}-\omega_{C}^{(t)}\right\rVert_{1}\leq 2\epsilon N$. Hence $I-(2\epsilon N+1)\leq\sum\limits_{i\in[I]}(1-\omega_{i}^{(t)})\leq 2\epsilon N$, which leads to $I\leq 4\epsilon N+1$. Therefore, according to Lemma B.15, with high probability, we have
$v^{\top}\left(\sum\limits_{i\in[I]\cap([N]\backslash C)}\omega_{i}^{(t)}Y_{i}Y_{i}^{\top}\right)v\leq c\gamma_{2}.$ (C.15)
@@
Therefore $\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{2}{3}\sum\limits_{i\in C}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{2(1-c)}{3}\gamma_{2}$. Note that by selecting the constant in (C.13) sufficiently large, we can let $c$ sufficiently small. Hence, we can assume $c\leq\frac{1}{4}$, by which we have $\sum\limits_{i\in[I]}\omega_{i}^{(t)}\tau_{i}^{(t)}\geq\frac{1}{2}\gamma_{2}$. Combine with (C.15), we find
@@
where the last step is from the definition of $\gamma_{2}$ (C.13). If we expand $v^{\top}Y_{i}$, this merely says
$\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]^{2}+\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}+2\sum\limits_{i\in[N]\backslash C}\omega_{i}\left[v^{\top}(Y_{i}-\nu)\right]\left\lVert\nu\right\rVert_{2}\leq cN.$ (C.16)
@@
Adding both sides with $\sum\limits_{i\in[N]\backslash C}\omega_{i}\left\lVert\nu\right\rVert_{2}^{2}$ and plugging (C.16) in, we have
@@
Now since $\left\lVert v\right\rVert_{2}=1$, we know that $\left[v^{\top}(Y_{i}-\nu)\right]^{2}$ is subject to a chi-square distribution with degree of freedom 1. By the standard argument for the sub-exponential random variables (see B.7), we know that $\sum\limits_{i\in[N]\backslash C}\left[v^{\top}(Y_{i}-\nu)\right]^{2}\leq N+c\sqrt{N}$ with (high) constant probability. On the other hand, by the definition of the algorithm we have $\left\lVert\omega_{[N]\backslash C}\right\rVert_{1}\geq\left\lVert\omega\right\rVert_{1}-\epsilon N\geq(1-3\epsilon)N$. Therefore, we have with (high) constant probability,
@@
Now deviding both sides by $\frac{1}{2}(1-3\epsilon)N$, we see that Lemma C.13 holds if we take $c=\frac{2(c_{1}+c_{2})}{1-3\epsilon}$
We are going to prove (iii) by contradiction. Suppose that at the end of the algorithm without rejection, we have $\left\lVert\sum\limits_{i\in S}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}>2\gamma_{2}\epsilon N$ for some $S\subset[N]$ with $\left|S\right|\leq\epsilon N$. This indicates that there exists a vector $v\in\mathbb{R}^{k^{\star}}$ with $\left\lVert v\right\rVert_{2}=1$ such that $\sum\limits_{i\in S}\sqrt{\omega_{i}}v^{\top}Y_{i}>\sqrt{2\gamma_{2}\epsilon N}$. Therefore, by the Cauchy–Schwarz inequality, we have $\sum\limits_{i\in S\cup C}\omega_{i}(v^{\top}Y_{i})^{2}\geq 2\gamma_{2}$. This implies that
$\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}+\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}\right\rVert_{2}\geq\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}Y_{i}Y_{i}^{\top}\right\rVert_{2}\geq 2\gamma_{2}.$ (C.17)
@@
$\leq\text{I}+\text{II}+\text{III}+\left|S\cup C\right|\left\lVert\nu\right\rVert_{2}^{2}.$
where I, II and III are defined as
$\text{I}=\left\lVert\sum\limits_{i=1}^{N}\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}-N\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2},\text{II}=\left\lVert\sum\limits_{i=1}^{N}(1-\omega_{i})\widetilde{Y}_{i}\widetilde{Y}_{i}^{\top}\right\rVert_{2},\text{III}=\left\lVert\sum\limits_{i\in S\cup C}\omega_{i}\widetilde{Y_{i}}\widetilde{Y_{i}}^{\top}\right\rVert$.
By Lemma C.13, without loss of generality we can assume $\left\lVert\nu\right\rVert_{2}\leq c_{0}$ for some constant $c_{0}$. For the last term, since $\left|S\cup C\right|\leq 2\epsilon N$, obviously it can be bounded by $c\gamma_{2}$. We are going to deal with the other three parts respectively.